### v1

@@ C.1 Dust density corresponding to Hi and CO emission
Figure 37: Two-dimensional histograms showing the distribution of $V_{\rm d}$ and distances for the data presented in Fig. 5.
@@
We also considered the potential bias introduced by the mean density, $n_{\rm H}\right>$, in a tile in the HOG results. It is expected that $V_{\rm d}$ should have higher values toward tiles with higher densities due to the larger number of significant density gradients for the same area. Yet, a morphological correlation should not be exclusively found toward high-$n_{\rm H}\right>$ tiles but distributed across $n_{\rm H}\right>$. This is what we found in the distribution of $V_{\rm d}$ as a function of $n_{\rm H}\right>$ reported in Fig. 38. We found that the highest $V_{\rm d}$ are not exclusively found in high-$n_{\rm H}\right>$ tiles, thus suggesting the HOG is finding a correlation that is not only related to the amount of dust in a tile but to the coincidence between its distribution and that of the line emission tracer in the plane of the sky.
@@
Figure 38: Two-dimensional histograms showing the distribution of $V_{\rm d}$ and $n_{\rm H}\right>$ for the data presented in Fig. 5.
@@
$\chi_{vd,l}=\frac{(V_{\rm d})_{vd,l}}{\left(2\sum_{ij}w_{ij,vd,l}\right)^{1/2}}.$ (10)
The quantity of $V_{d,{\rm max}}$ represents the maximum value of Eq. (2), which corresponds to all gradient pairs being parallel, $\theta_{ij,lm}$$=$ 0, or the comparison of identical images. In the particular case $w_{ij,lm}$$=$ 1, $\left(V_{d,{\rm max}}\right)_{lm}$$=$$(2N_{lm})^{1/2}$, where $N_{lm}$ is the number of relative orientation angles between gradients in velocity channel $l$ and distance channel $m$. The values of $\chi$ are bound to the range $-1$ and $1$ and roughly correspond to the percentage of the gradient pairs representing the parallel or antiparallel direction, for $V_{\rm d}$$>$ 0 and $V_{\rm d}$$<$ 0, respectively.
@@
Figure 39: Same as Fig. 5, but for the normalized projected Rayleigh statistic, $\chi$, Eq. 10.
C.2 Assignment of $v_{\rm LOS}$ to dust channels
Once we identified a significant correlation between the 3D dust and the line emission, we assigned a velocity to each distance channel as follows. We ignored the distance-$v_{\rm LOS}$ pairs for which $V_{\rm d}$$<$ 2.87 or $V_{\rm d}$$<$$\sigma_{V_{\rm d}}$, where $\sigma_{V_{\rm d}}$ is the standard deviation of the $V_{\rm d}$ values found for the 12 samples of the 3D dust reconstruction and the $N_{\rm MC}$ samples of the line emission described in App. A. For each distance channel, we selected the $v_{\rm LOS}$ corresponding to the highest $V_{\rm d}$ in each tracer, as illustrated by the white markers in Fig. 25. Generally, the highest $V_{\rm d}$ values are clustered around a single central $v_{\rm LOS}$, as expected from the correlation in velocity introduced by the emission line width. There are, however, a few cases where two line-of-sight velocities separated by more than 10 km s-1 show singularly high $V_{\rm d}$ values. This case is most likely the result of two dominant velocity components in the physical volume comprised of one tile, which is exciting but complicates the large-scale study considered in this paper. For the sake of simplicity, we assigned to a distance channel only the $v_{\rm LOS}$ corresponding to the highest $V_{\rm d}$ and reserved other dynamical effects revealed by the HOG analysis to a follow-up publication.
We evaluated the convergence of this $v_{\rm LOS}$ assignment by considering the three Galactic plane segmentations described in App. A.3.2. Figure 40 shows the distribution of the $v_{\rm LOS}$ assigned to the dust tiles using the Hi and CO emission. We found that the assigned $v_{\rm LOS}$ values have a global distribution primarily unaffected by the sky segmentation. The discrepancies between the three $v_{\rm LOS}$ distributions indicate the suppression of some velocity features toward particular lines of sight, which are worth exploring in a focused study of specific regions. The regularity in the global distribution indicates that the average physical quantities derived from this $v_{\rm LOS}$ do not critically depend on our division of the sky band $|b|$$<$ 5.
Figure 41: Histogram of the effective densities assigned to Hi and CO across distance channels.
We computed the effective nucleon density for gas tracer $X$ in the distance channel $m$ as
$n^{X}_{m}=\frac{\sum_{ij}w_{ij}n_{ij,m}A_{ij}\Delta z_{m}}{\sum_{ij}A_{ij}\Delta z_{m}}=\frac{\sum_{ij}w_{ij}n_{ij,m}A_{ij}}{\sum_{ij}A_{ij}},$ (11)
where $(n_{H})_{ij,m}$ is the nucleon density derived from the 3D dust reconstruction, $A_{ij}$ is the pixel area, $\Delta z_{m}$ is the distance channel width, and $w_{ij}$ is a statistical weight set to 1.0 if the pixel is within a block with $V_{\rm d}$$>$ 0 and 0.0 otherwise. Figure 41 shows the distribution of effective densities for the two gas tracers. The mean effective densities are roughly 0.5 and 0.25 times the mean density from the 3D dust reconstruction for Hi and CO, respectively. The lower effective CO density is product of the lower filling factor of that line emission in the maps and roughly reflects the larger mass of the atomic gas in the Milky Way; between 9.4 and 10.5 $\times$$10^{9}$ M of atomic gas (Wang et al. 2020a) and 1.0 $\pm$0.3 M of H2 (Heyer & Dame 2015).
@@ C.2 Assignment of $v_{\rm LOS}$ to dust channels
Appendix D Comparison with parallax and proper motions of star-forming regions
@@
Appendix E Distance-LOS velocity correlation toward the Galactic center and anticenter
The assumption of circular motions around the Galactic center to derive distances from the observed emission across $v_{\rm LOS}$ is particularly catastrophic in directions where the LOS component of the velocity is small, for example, toward the Galactic center and anticenter (see, for example, Hunter et al. 2024, and references their). Therefore, we report the HOG-derived distance-$v_{\rm LOS}$ mapping for four reference regions in Fig. 43.
Toward the Galactic center, the analysis shows that the highest $V_{\rm d}$ are distributed in dust parcels located at $d$$\approx$ 950 and 1125 pc However, these clouds appear distributed over a broad range of $v_{\rm LOS}$  both in Hi and CO. Toward the Galactic anticenter, the highest $V_{\rm d}$ values appear distributed across multiple distances in more restricted $v_{\rm LOS}$ ranges. For CO the highest $V_{\rm d}$ appears around $v_{\rm LOS}$$\approx$ 10 and $-10$ km s-1 across distances between 69 and 600 pc, although the radial velocity expectation is close to zero. There is a significant offset between the Hi and CO with the highest $V_{\rm d}$ for the same distance channels, showing amplitudes that are comparable to the global velocity offsets reported in Fig. 22.

### v2

@@ Appendix C Effects of fixed angular resolution and distance
Figure 41: Two-dimensional histograms showing the distribution of $V_{\rm d}$ and distances for the data presented in Fig. 5. The solid white lines correspond to the 50th and 99th percentile of the $\max(V_{\rm d})$ values in each distance bin.
@@
We also considered the potential bias introduced by the mean density, $n_{\rm H}\right>$, in a tile in the HOG results. It is expected that $\max(V_{\rm d})$ should have higher values toward tiles with higher densities due to the larger number of significant density gradients for the same area. Yet, a morphological correlation should not be exclusively found toward high-$n_{\rm H}\right>$ tiles but distributed across $n_{\rm H}\right>$. This is what we found in the distribution of $V_{\rm d}$ as a function of $n_{\rm H}\right>$ reported in Fig. 42. We found that the highest $V_{\rm d}$ are not exclusively found in high-$n_{\rm H}\right>$ tiles, thus suggesting the HOG is finding a correlation that is not only related to the amount of dust in a tile but to the coincidence between its distribution and that of the line emission tracer in the plane of the sky.
@@
Figure 42: Two-dimensional histograms showing the distribution of $V_{\rm d}$ and $n_{\rm H}\right>$ for the data presented in Fig. 5. The solid white lines correspond to the 50th percentile of the $\max(V_{\rm d})$ values in each density bin.
@@
$\chi_{kpq}=\frac{(V_{\rm d})_{kpq}}{(V_{\rm d,max})_{kpq}}=\frac{(V_{\rm d})_{kpq}}{\left(2\sum_{ij}w_{ijkpq}\right)^{1/2}}.$ (10)
The quantity of $V_{d,{\rm max}}$ represents the maximum value of Eq. (2), which corresponds to all gradient pairs being parallel, $\theta_{ijkpq}$$=$ 0, or the comparison of identical images. In the particular case $w_{ijkpq}$$=$ 1, $\left(V_{d,{\rm max}}\right)_{kpq}$$=$$(2N_{kpq})^{1/2}$, where $N_{kpq}$ is the number of relative orientation angles between gradients in distance channel $p$ and velocity channel $q$. The values of $\chi$ are bound to the range $-1$ and $1$ and roughly correspond to the percentage of the gradient pairs representing the parallel or antiparallel direction, for $V_{\rm d}$$>$ 0 and $V_{\rm d}$$<$ 0, respectively.
Figure 43 shows the maximum $\chi$ across $v_{\rm LOS}$ channels for each distance channel in our reconstruction. This quantity can be considered an analog of $\max(V_{\rm d})$ in Fig. 5, but marginalizing over the difference in the number of independent gradients across tiles. Visual comparison between Fig. 5 and Fig. 43 indicates that the $\max(V_{\rm d})$ and $\max(\chi)$ distributions are very similar in Hi, as expected from the homogeneous number of independent gradients across distance and $v_{\rm LOS}$ channels. However, in CO, some portions of the $b$$\leq$ 5range are not uniform, producing a dissimilar number of independent gradients across Galactic longitudes and $v_{\rm LOS}$. This effect is evident in the differences between $\max(V_{\rm d})$ and $\max(\chi)$ for that tracer toward the third Galactic quadrant and other low CO emission regions. In those regions, the normalization implied in Eq. 10 may emphasize a morphological correlation represented by just a few independent gradients. Thus, we settle for $V_{\rm d}$ as a morphological metric rather than $\chi$.
@@
Figure 43: Same as Fig. 5, but for the normalized projected Rayleigh statistic, $\chi$, Eq. (10).
Appendix D The HOG method in synthetic observations
We conducted a HOG study of the synthetic Hi and CO observations of a simulated MC in the Simulating the life-Cycle of molecular clouds (SILCC) Zoom project (Seifried et al. 2017, 2020). It corresponds to a 125-pc-side cube extracted for the kiloparsec-sized stratified boxes in the SILCC project (Walch et al. 2015; Girichidis et al. 2016), modeled at 0.06-pc resolution using adaptive mesh refinement in combination with a chemical network to follow heating, cooling and the chemistry of Hi and the formation of H2 and CO including (self-) shielding. We used their MC1-MHD model, one of the four clouds studied through Hi synthetic observations in Seifried et al. (2022). We refer to the aforementioned publications for further details on the simulations and synthetic observations. .
We used the synthetic Hi and 12CO(1-0) from the SILCC-Zoom MC1-MHD simulation to mimic observations comparable to those employed in this paper. We placed the simulated cube at a distance of 709 pc parsec to cover the range $l$$<$±$\pm 5$. We smoothed each observation using a 2D-Gaussian point spread function matching the corresponding beam and used an identical grid to that employed in the analysis of observational data, as described in Sec. 2. As a proxy for the 3D dust PPD cube, we used the 3D density cube from the simulation but smoothed employing a 2D-Gaussian kernel to match the angular resolution of the Edenhofer et al. (2024) reconstruction and projected it into a regular distance grid with a channel width of 0.24 pc. The resulting maps are presented in Fig. 44.
Figure 44: Projected synthetic 12CO(1-0) and Hi line emission and 3D dust reconstruction from the SILCC Zoom MC1-MHD simulation placed at 709 pc from the LSR.
We applied the HOG analysis to the MC1-MHD synthetic observations using the same parameters described in Sec. 3. The results, shown in Fig. 45, confirm the capabilities of HOG to identify the line emission from an object in the PPD density cube. We employed this result to evaluate the density and $v_{\rm LOS}$ reconstructions presented in Sec. 5.
Figure 45: Same as Fig. 4, but for the synthetic observations of the SILCC-Zoom MC1-MHD simulation.
Figure 46 shows a comparison of the densities across LOS distance channels obtained directly from the simulation and reconstructed with the HOG using Eq. (3). The reconstructed densities for Hi are roughly within a factor of two from their values in the simulation and recover the global density profile along the LOS. The values reconstructed with the HOG are smaller than the actual values, which is an expected consequence of estimating the densities using block averaging. Reconstructions and line emission observations with higher angular resolutions enable finer segmentation and better estimates. For CO, the HOG method tends to overestimate the density, possibly by smearing the concentrated emission along the block area. However, this comparison is limited by the amount of extended, diffuse CO emission in this model, which is still a subject of active discussion (e.g., Glover & Clark 2012; Levrier et al. 2012; Godard et al. 2023; Beitia-Antero et al. 2024)
Regarding $v_{\rm LOS}$ reconstruction, the HOG method correctly reproduces the global LOS motion in the MC1-MHD simulation, which is 0 km s-1. The width of the $V_{\rm d}$ signal in the $v_{\rm LOS}$ axis of Fig. 45 corresponds to the emission linewidth coming from this isolated object (Seifried et al. 2022). Excursions from the central $v_{\rm LOS}$ are most likely related to the internal motions of the cloud and its interaction with the environment. Further identification of the internal MC dynamics using the HOG is beyond the scope of this work. However, this numerical experiment indicates that the recovered $v_{\rm LOS}$ dispersion is related to the Hi and CO emission linewidths.
Figure 46: Comparison between the total, Hi, and CO densities across distance slices obtained directly from the simulation cube and reconstructed with HOG using Eq. (3).
@@ Appendix D The HOG method in synthetic observations
Appendix E Comparison with parallax and proper motions of star-forming regions
@@
Appendix F Distance-LOS velocity correlation toward the Galactic center and anticenter
The assumption of circular motions around the Galactic center to derive distances from the observed emission across $v_{\rm LOS}$ is particularly catastrophic in directions where the LOS component of the velocity is small, for example, toward the Galactic center and anticenter (see, for example, Hunter et al. 2024, and references therein). Therefore, we report the HOG-derived distance-$v_{\rm LOS}$ mapping for four reference regions in Fig. 48.
Toward the Galactic center, the analysis shows that the highest $V_{\rm d}$ are distributed in dust parcels located at $d$$\approx$ 950 and 1125 pc It is apparent that the signal from the dust clouds in that distance range is smeared across a broad $v_{\rm LOS}$ range. This is unlikely the result of the cloud dynamics but rather a manifestation of the mapping of one density structure into multiple velocities in line emission due to the large velocity gradients along the LOS (see, for example, Beaumont et al. 2013). The presence of substantial $v_{\rm LOS}$ gradients and the high densities sampled across this LOS also implies that a significant portion of the line emission may originate from distances beyond the range of the reconstruction. However, testing this hypothesis requires deeper LOS 3D dust reconstructions, which are not yet available.
Toward the Galactic anticenter, the highest $V_{\rm d}$ values appear distributed across multiple distances in a limited $v_{\rm LOS}$ range. This is an archetypal example of the effect known as “velocity crowding”, which is the superposition of multiple PPD structures into a few channels in PPV (see, for example, Ballesteros-Paredes & Mac Low 2002). For CO, the highest $V_{\rm d}$ appears around $v_{\rm LOS}$$\approx$ 10 and $-10$ km s-1 across distances between 69 and 600 pc, although the radial velocity expectation is close to zero. There are also considerable offsets from the expected $v_{\rm LOS}$ in Hi. For both gas tracers, the amplitude of the departure from the expected $v_{\rm LOS}$ is comparable to the global velocity offsets reported in Fig. 22.